# Option-aware Temporally Abstracted Value for Offline Goal-Conditioned Reinforcement Learning

### v1

@@ 5 Option-aware Temporally Abstracted (OTA) Value
In this section, we propose a simple solution for learning $V^{h}(s,g)$ by leveraging options [46] to reduce the horizon length. Option, also known as macro actions, is a sequences of primitive actions that enable temporal abstraction. In our offline RL setting, an option starting from the state $s_{t}$ corresponds to a sequence of $n$ actions $(a_{t},a_{t+1},\ldots,a_{t+n-1})$, which are extracted from trajectories in the offline dataset. By using temporally extended actions in planning, we reduce the effective horizon length, referring to the number of planning steps, to approximately $d^{\star}(s_{t},g)/n$. Therefore, to ensure that the high-level value $V^{h}$ is suitable for long-term planning, we modify the reward and target value in Equation (1) to be option-aware.
More specifically, for a given abstraction factor $n$ and goal $g$, we define an option $\Omega_{n,g}=({\mathcal{I}},\mu,\beta_{n,g})$, where ${\mathcal{I}}={\mathcal{S}}$ is the initiation set, $\mu$ is the behavior policy used to collect the offline dataset $\mathcal{D}$}, and $\beta_{n,g}$ is a timeout-based termination condition that ends the option after $n$ steps or upon reaching $g$. Let $s^{\prime}(\Omega_{n,g},s_{t})$ denote the state resulting from executing $\Omega_{n,g}$ at state $s_{t}$, which is either $s_{t+n}$ or $g$. For brevity, we denote $s^{\prime}(\Omega_{n,g},s)$ as $s^{\Omega}$. Then, we reformulate the value learning objective in Equation (1) into an option-aware variant as follows:
@@
We argue that the high-level value function $V^{h}_{\text{OTA}}$ would effectively address the value order inconsistency due to the following reason: Using a $1$-step target for value learning tends to be more sensitive to noise, particularly in long-horizon tasks. In contrast, employing an option-aware target mitigates the effects of noise and empirically leads to more order-consistent value estimates. The overall framework for learning $V^{h}_{\text{OTA}}$ is illustrated in Figure 3.
@@ 6.1 Experiment Setup
Tasks. We use OGBench [33], a recently proposed offline GCRL benchmark designed for realistic environments, long-horizon scenarios, and multi-goal evaluation. The Maze environment consists of long-horizon navigation tasks that evaluate whether the agent can reach a specified goal position from a given initial state. The Maze environments are categorized by agent type (PointMaze, AntMaze, and HumanoidMaze), maze size (medium, large, and giant), and the type of trajectories in the dataset (navigate, stitch, and explore). The Maze environments are well suited to evaluating performance in long-horizon settings. For example, the HumanoidMaze-giant environment has a maximum episode length of 4000 steps. The Visual-cube and Visual-scene environments focus on visual robotic manipulation tasks. In Visual-cube, the task involves manipulating and stacking cube blocks to reach a specified goal configuration. This environment is categorized by the number of cubes: single, double, and triple. In contrast, Visual-scene requires the agent to control everyday objects like windows, drawers, or two-button locks in a specific sequence. Both visual environments use high-dimensional, pixel-based observations with $64\times 64\times 3$ RGB images. The robotic manipulation environments have shorter episode lengths ($250$ to $1000$ steps) compared to the Maze tasks. These robotic manipulation environments are a strong benchmark for evaluating the performance of an algorithm on high-dimensional visual inputs. A detailed description of the environments is provided in Appendix B.1.
Baselines. For brevity, we will refer to the policy that utilizes the high-level policy learned with the OTA value as OTA. We compare OTA against six representative offline GCRL methods included in OGBench. Goal-conditioned behavioral cloning (GCBC[11] is a simple behavior cloning method that directly imitates actions from the dataset conditioned on the goal. Goal-conditioned implicit V-learning (GCIVL) and goal-conditioned implicit Q-learning (GCIQL[22, 34] estimate the goal-conditioned optimal value function using IQL-based expectile regression, and extract policies using AWR [37] and behavior-regularized deep deterministic policy gradient (DDPG+BC) [10], respectively. Quasimetric RL (QRL[51] learns a value function that estimates the undiscounted temporal distance between state and goal via quasimetric learning and trains a policy using DDPG+BC. Contrastive RL (CRL[9] approximates the Q-function via contrastive learning between state-action pairs and future states from the same trajectory, and trains the policy using DDPG+BC. HIQL [34] extends GCIVL with a hierarchical policy, as detailed in Section 3.
@@ 6.4 Effect of Varying Abstraction Factor $n$
The training of the value function $V^{h}_{\text{OTA}}$ depends on the abstraction factor $n$, which determines the degree of temporal abstraction. Figure 6(a-c) illustrates how the value function changes as $n$ is varied across $1,2,3,5,10$, and $20$ in Equation 4, while keeping the optimal trajectory and goal fixed for each dataset. As shown in Figure 6(b,c), for long-horizon trajectories (i.e., those exceeding a length of 1500), the absolute scale of the value function increases with larger $n$. This trend arises since the option termination condition introduces a reward of 1-1 every $n$ steps, which effectively compresses the value range as $n$ increases.
@@
6.5 Effect of Scaling the Discount Factor $\gamma$
Table 1: Average success rate and order consistency ratio. Simply increasing the discount factor in HIQL is insufficient to achieve the performance improvements of OTA.
@@ 6.5 Effect of Scaling the Discount Factor $\gamma$
HIQL $(\gamma)$ HIQL $(\gamma^{1/n})$ OTA HIQL $(\gamma)$ HIQL $(\gamma^{1/n})$ OTA
AntMaze-large-explore 4 $\scriptstyle\pm 5$} 3 $\scriptstyle\pm 3$} 75 $\scriptstyle\pm 16$} 0.75 $\scriptstyle\pm 0.01$} 0.76 $\scriptstyle\pm 0.02$} 0.97 $\scriptstyle\pm 0.01$}
AntMaze-giant-stitch 2 $\scriptstyle\pm 2$} 0 $\scriptstyle\pm 0$} 37 $\scriptstyle\pm 6$} 0.91 $\scriptstyle\pm 0.01$} 0.79 $\scriptstyle\pm 0.02$} 0.94 $\scriptstyle\pm 0.01$}
HumanoidMaze-large-stitch 12 $\scriptstyle\pm 4$} 22 $\scriptstyle\pm 3$} 57 $\scriptstyle\pm 3$} 0.76 $\scriptstyle\pm 0.01$} 0.75 $\scriptstyle\pm 0.02$} 0.89 $\scriptstyle\pm 0.03$}
HumanoidMaze-giant-stitch 28 $\scriptstyle\pm 3$} 2 $\scriptstyle\pm 1$} 79 $\scriptstyle\pm 3$} 0.71 $\scriptstyle\pm 0.01$} 0.72 $\scriptstyle\pm 0.01$} 0.94 $\scriptstyle\pm 0.01$}
In the original HIQL, the high-level value function $V^{h}$ is discounted by $\gamma$ at every step. In contrast, the OTA value function $V^{h}_{\text{OTA}}$ applies discounting every $n$ steps, leading to a more increased optimal value function $V^{\star}(s_{t},g)=-(1-\gamma^{d^{\star}(s_{t},g)/n})/(1-\gamma)$. A straightforward approach to mimic a temporally abstracted optimal value function is to increase the discounting factor $\gamma$ in Equation (1). Specifically, we modify the standard discount factor by using $\gamma^{1/n}$ instead of $\gamma$ for training the high-level value function of HIQL. We compare standard HIQL, which uses the original discount factor $\gamma$ (denoted as HIQL ($\gamma$)), with a variant learned using a modified discount factor $\gamma^{1/n}$ (denoted as HIQL ($\gamma^{1/n}$)).
We evaluate both the success rate and the order consistency ratio $r^{c}$ across four datasets. For OTA and HIQL ($\gamma^{1/n}$), we use $n=15$ for AntMaze and $n=20$ for HumanoidMaze. To compute $r^{c}$, we collect 5 trajectories per dataset and report the average consistency ratio (see Appendix B.2.3 for details of the collected trajectories). Table 1 shows that HIQL ($\gamma^{1/n}$) fails to outperform standard HIQL in either success rate or $r^{c}$ in most cases, whereas OTA achieves significant gains in long-horizon tasks such as HumanoidMaze-giant-stitch. These results highlight that simply altering the discounting factor is insufficient and that temporal abstraction is crucial for effective value learning in long-horizon environments.
@@ 6.6 Scalability Comparison of TD-Based OTA and QRL
Here, we demonstrate that OTA, which leverages a TD-based IQL loss, scales effectively with increasing state and action dimensionality. As discussed in Section 4.2, conventional TD methods use a discount factor, leading to exponential decay of the advantage signal in long-horizon settings. This decay weakens the advantage signal in long-horizon settings and can hinder high-level policy learning in the absence of OTA.
To address the limitation, we consider alternative value learning approaches that do not rely on a discount factor. One such approach is QRL, which learns undiscounted temporal distances between states through quasimetric learning (see Appendix C for more details). However, QRL relies on min-max optimization, which becomes computationally challenging in high-dimensional state spaces. As shown in Table 2, QRL achieves significantly lower success rates on complex tasks such as HumanoidMaze and Visual-scene. These results highlight the scalability and the practical advantages of our TD-based OTA, particularly in environments with high-dimensional state spaces.
@@ 7 Conclusion
In this paper, we investigated the limitations of the hierarchical offline GCRL method HIQL, particularly in long-horizon tasks. Our analysis revealed that the main performance bottleneck lay in the high-level policy, which suffers from inaccurate value estimates when the state and goal are far apart. To address this challenge, we proposed OTA, a method that incorporates temporal abstraction into IQL-based value learning by leveraging the concept of options. Our experiments on challenging long-horizon goal-reaching tasks demonstrated that high-level policies learned with OTA achieved significant performance gains in long-term planning. We believe that the simplicity and effectiveness of OTA present a promising direction for long-horizon offline GCRL in real-world applications.
@@ B.2.1 Hyperparameters
B.2.2 Training and Evaluation detail
@@ B.2.3 Collected Optimal trajectories
The collected optimal trajectories for AntMaze and HumanoidMaze are shown in Figures 8 and 9, respectively. Order consistency, as reported in Table 1, is evaluated based on the five trajectories illustrated in these figures and averaged over $8$ random seeds. The optimal trajectories used for value visualizations in Figures 5 and 6 are as follows:

### v2

@@ 5 Option-aware Temporally Abstracted (OTA) Value
In this section, we propose a straightforward solution for learning $V^{h}(s,g)$ by leveraging options (option)Sutton1999 to reduce the horizon length. An option can be regarded as a temporally extended sequence of primitive actions that enable temporal abstraction. In our offline RL setting, an option starting from the state $s_{t}$ corresponds to a sequence of $n$ actions $(a_{t},a_{t+1},\ldots,a_{t+n-1})$ extracted from trajectories in the offline dataset. By using temporally extended actions in planning, we reduce the effective horizon length, referring to the number of planning steps, to approximately $d^{\star}(s_{t},g)/n$. Therefore, to ensure that the high-level value $V^{h}$ is suitable for long-term planning, we modify the reward and target value in Equation (1) to be option-aware.
Specifically, for a given abstraction factor $n$ and goal $g$, we define an option $\Omega_{n,g}=({\mathcal{I}},\mu,\beta_{n,g})$, where ${\mathcal{I}}={\mathcal{S}}$ is the initiation set, $\mu$ is the behavior policy used to collect the offline dataset $\mathcal{D}$}, and $\beta_{n,g}$ is a timeout-based termination condition that ends the option after $n$ steps or upon reaching $g$. Let $s^{\prime}(\Omega_{n,g},s_{t})$ denote the state resulting from executing $\Omega_{n,g}$ at state $s_{t}$, which is either $s_{t+n}$ or $g$. For brevity, we denote $s^{\prime}(\Omega_{n,g},s)$ as $s^{\Omega}$. Then, we reformulate the value learning objective in Equation (1) into an option-aware variant:
@@
We argue that the high-level value function $V^{h}_{\text{OTA}}$ would effectively address the value order inconsistency. Using a $1$-step target for value learning tends to be more sensitive to noise, especially in long-horizon tasks, whereas an option-aware target mitigates noise and empirically produces more order-consistent value estimates. The overall framework for learning $V^{h}_{\text{OTA}}$ is illustrated in Figure 3.
Connection to $n$-step TD learning. The target used in $n$-step TD learning and that in OTA value learning are closely related, as both primarily rely on the $n$-step forward value. [^6] The TD target used in $n$-step TD learning is $\sum_{i=0}^{n-1}-\gamma^{i}\cdot\mathbf{1}(s_{t+i}\neq g)+\gamma^{n}V(s_{t+n},g)$. However, the key distinction between $n$-step TD and OTA lies in the choice of the discount factor $\gamma$, which controls how information decays during the TD update. Standard $n$-step TD learning typically uses the same $\gamma$ as in 1-step TD, causing the discount factor applied to the $n$-step target to decay exponentially with $n$. In contrast, the discount factor in the OTA target is independent of $n$. This excessive decay in the standard $n$-step target hinders the order-consistent value learning, indicating that a direct extension from the $n$-step target to the OTA target is not straightforward. Instead, temporal abstraction through the option framework provides a natural explanation for the insights presented in Section 4.2. As shown empirically in 6.5, with respect to value order consistency, standard $n$-step TD learning offers no advantage over $1$-step TD learning.
@@ 6.1 Experiment Setup
Tasks. We use OGBench (OGBench)Park2025 , a recently proposed offline GCRL benchmark designed for realistic environments, long-horizon scenarios, and multi-goal evaluation. The Maze environment consists of long-horizon navigation tasks that evaluate whether the agent can reach a specified goal position from a given initial state. The Maze environments are categorized by agent type (PointMaze, AntMaze, and HumanoidMaze), maze size (medium, large, and giant), and the type of trajectories in the dataset (navigate, stitch, and explore). The Maze environments are well suited to evaluating performance in long-horizon settings. For example, the HumanoidMaze-giant environment has a maximum episode length of 4000 steps.
The Visual-cube and Visual-scene environments focus on visual robotic manipulation tasks. In Visual-cube, the task involves manipulating and stacking cube blocks to reach a specified goal configuration. This environment is categorized by the number of cubes: single, double, and triple. In contrast, Visual-scene requires the agent to control everyday objects like windows, drawers, or two-button locks in a specific sequence. Both visual environments use high-dimensional, pixel-based observations with $64\times 64\times 3$ RGB images. The robotic manipulation environments have shorter episode lengths ($250$ to $1000$ steps) compared to the Maze environments. These robotic environments are a strong benchmark for evaluating the performance of an algorithm on high-dimensional visual inputs. A detailed description of the environments is provided in Appendix B.1.
Baselines. For brevity, we will refer to the policy that utilizes the high-level policy learned with the OTA value as OTA. We compare OTA against six representative offline GCRL methods included in OGBench. Goal-conditioned behavioral cloning (GCBC(GCSL)Ghosh2019 is a simple behavior cloning method that directly imitates actions from the dataset conditioned on the goal. Goal-conditioned implicit V-learning (GCIVL) and goal-conditioned implicit Q-learning (GCIQL(IQL)Kostrikov ; (HIQL)Park2023 estimate the goal-conditioned optimal value function using IQL-based expectile regression, and extract policies using AWR (AWR)Peng2019 and behavior-regularized deep deterministic policy gradient (DDPG+BC) (DDPG+BC)fujimoto2021 , respectively. Quasimetric RL $(QRL$)Wang2023 learns a value function that estimates the undiscounted temporal distance between state and goal via quasimetric learning and trains a policy using DDPG+BC. Contrastive RL $(CRL$)Eysenbach2022 approximates the Q-function via contrastive learning between state-action pairs and future states from the same trajectory, and trains the policy using DDPG+BC. HIQL (HIQL)Park2023 extends GCIVL with a hierarchical policy, as detailed in Section 3.
@@ 6.4 Effect of Varying Abstraction Factor $n$
Learning the value function $V^{h}_{\text{OTA}}$ depends on the abstraction factor $n$, which determines the degree of temporal abstraction. Figure 6(a-c) illustrates how the value function changes as $n$ is varied across $1,2,3,5,10$, and $20$ in Equation 4, while keeping the optimal trajectory and goal fixed for each dataset. As shown in Figure 6(b,c), for long-horizon trajectories (i.e., those exceeding a length of 1500), the absolute scale of the value function increases with larger $n$. This trend arises since the option termination condition introduces a reward of 1-1 every $n$ steps, which effectively compresses the value range as $n$ increases.
@@
6.5 Impact of $n$-Step TD and Increasing the Discount Factor $\gamma$
Table 1: Average success rate and order consistency ratio. Simply using $n$-step TD learning or increasing the discount factor in HIQL is insufficient to achieve the performance improvements of OTA. Here, the dataset ALE refers to AntMaze-large-explore, AGS to AntMaze-giant-stitch, HLS to HumanoidMaze-large-stitch, and HGS to HumanoidMaze-giant-stitch.
@@ 6.5 Impact of $n$-Step TD and Increasing the Discount Factor $\gamma$
HIQL OTA HIQL OTA
1-step, $\gamma$ $n$-step, $\gamma$ 1-step, $\gamma^{1/n}$ 1-step, $\gamma$ $n$-step, $\gamma$ 1-step, $\gamma^{1/n}$
ALE 4 $\scriptstyle\pm 5$} 0 $\scriptstyle\pm 0$} 3 $\scriptstyle\pm 3$} 75 $\scriptstyle\pm 16$} 0.75 $\scriptstyle\pm 0.01$} 0.77 $\scriptstyle\pm 0.01$} 0.76 $\scriptstyle\pm 0.02$} 0.97 $\scriptstyle\pm 0.01$}
AGS 2 $\scriptstyle\pm 2$} 0 $\scriptstyle\pm 0$} 0 $\scriptstyle\pm 0$} 37 $\scriptstyle\pm 6$} 0.91 $\scriptstyle\pm 0.01$} 0.84 $\scriptstyle\pm 0.02$} 0.79 $\scriptstyle\pm 0.02$} 0.94 $\scriptstyle\pm 0.01$}
HLS 12 $\scriptstyle\pm 4$} 50 $\scriptstyle\pm 4$} 22 $\scriptstyle\pm 3$} 57 $\scriptstyle\pm 3$} 0.76 $\scriptstyle\pm 0.01$} 0.76 $\scriptstyle\pm 0.00$} 0.75 $\scriptstyle\pm 0.02$} 0.89 $\scriptstyle\pm 0.03$}
HGS 28 $\scriptstyle\pm 3$} 2 $\scriptstyle\pm 2$} 2 $\scriptstyle\pm 1$} 79 $\scriptstyle\pm 3$} 0.71 $\scriptstyle\pm 0.01$} 0.72 $\scriptstyle\pm 0.00$} 0.72 $\scriptstyle\pm 0.01$} 0.94 $\scriptstyle\pm 0.01$}
In the original HIQL, the high-level value function $V^{h}$ is discounted by $\gamma$ at every step. In contrast, the OTA value function $V^{h}_{\text{OTA}}$ applies discounting only every $n$ steps. To investigate the source of the effectiveness of OTA, we modify the value learning approach of standard HIQL in two ways: (1) using $n$-step TD learning, and (2) increasing $\gamma$. We evaluate both the success rate and the order consistency ratio $r^{c}$ across four datasets. In Table 1, we set $n=15$ for AntMaze and $n=20$ for HumanoidMaze. To compute $r^{c}$, we collect 5 trajectories per dataset and report the average consistency ratio (see Appendix B.2.3 for details of the collected trajectories).
The first variant uses the original $\gamma$ with $n$-step TD learning in HIQL, denoted as HIQL($n$-step, $\gamma$). Table 1 shows that this approach yields almost no improvement in $r^{c}$, and the success rates also show little gain except for HumanoidMaze-large-stitch. These results indicate that $n$-step TD targets still suffer from value function estimation errors when the discount factor remains unchanged.
The second variant keeps 1-step TD learning but modifies the discount factor to $\gamma^{1/n}$. Under OTA training, the optimal value function becomes $V^{\star}(s_{t},g)=-(1-\gamma^{d^{\star}(s_{t},g)/n})/(1-\gamma)$. Therefore, to approximate this temporally abstracted optimal value function, a possible approach is to increase the discount factor $\gamma$ to $\gamma^{1/n}$ in Equation (1). However, Table 1 shows that simply increasing $\gamma$ fails to outperform standard HIQL in either success rate or $r^{c}$. In contrast, OTA achieves significant gains in long-horizon tasks such as HumanoidMaze-giant-stitch. The experiments demonstrate that simply adjusting the discounting factor alone is insufficient, and the temporal abstraction is crucial for effective value learning in long-horizon tasks.
Our analysis further suggests that $n$-step TD learning could potentially be improved by carefully adjusting $\gamma$ for each $n$. However, this would introduce additional complexity in hyperparameter selection. In contrast, OTA fixes $\gamma$ regardless of $n$, which makes the approach much simpler.
@@ 6.6 Scalability Comparison of TD-Based OTA and QRL
Here, we demonstrate that OTA, which leverages a TD-based IQL loss, scales effectively with increasing state and action dimensionality. As discussed in Section 4.2, conventional TD methods rely on a discount factor, which causes the advantage estimate to decay exponentially over long horizons. To avoid this issue, we explore alternative value learning approaches that do not depend on a discount factor.
@@
In particular, we consider QRL, which learns undiscounted temporal distances between states through quasimetric learning (see Appendix C for more details). However, QRL relies on min-max optimization, which becomes computationally challenging in high-dimensional state spaces. As shown in Table 2, QRL achieves significantly lower success rates on complex tasks such as HumanoidMaze and Visual-scene. These results highlight the scalability and the practical advantages of our TD-based OTA, particularly in environments with high-dimensional state spaces.
@@ 7 Conclusion
In this paper, we investigated the limitations of the hierarchical offline GCRL method HIQL, particularly in long-horizon tasks. Our analysis revealed that the main performance bottleneck lies in the high-level policy, which suffers from inaccurate value estimates when the state-goal distance is large. To address this challenge, we proposed OTA, a method that incorporates temporal abstraction into IQL-based value learning through the concept of options. Experiments on challenging long-horizon goal-reaching tasks show that OTA enables high-level policies to achieve substantial performance improvements in long-term planning. The simplicity and effectiveness of OTA highlight its potential for advancing long-horizon offline GCRL in real-world applications.
Acknowledgments and Disclosure of Funding
This work was supported in part by National Research Foundation of Korea (NRF) grant [No. 2021R1A2C2007884, No. RS-2025-02263628], the Institute of Information & communications Technology Planning & Evaluation (IITP) grants [RS-2021-II212068, RS-2022-II220113, RS-2022-II220959, RS-2021-II211343], and the BK21 FOUR Education and Research Program for Future ICT Pioneers (Seoul National University), funded by the Korean government (MSIT).
@@ B.2.1 Hyperparameters
B.2.2 Training and Evaluation Details
@@ B.2.3 Collected Optimal Trajectories
The collected optimal trajectories for AntMaze and HumanoidMaze are shown in Figures 8 and 9, respectively. Order consistency, as reported in Table 1, is evaluated based on the five trajectories illustrated in these figures and averaged over $8$ random seeds. During value estimation, we apply a moving average with an appropriate temporal window size to smooth out short-term fluctuations and obtain stable value estimates. The optimal trajectories used for value visualizations in Figures 5 and 6 are as follows:
@@ D.1 Per-environment Results
D.2 Performance under Unified Hyperparameters
We report additional results where OTA is trained with the same hyperparameters as HIQL, except for the temporal abstraction factor $n$. The experiments are conducted on complex maze environments.
As shown in Table 7, OTA consistently outperforms HIQL even under identical hyperparameter settings. This result indicates that incorporating temporal abstraction alone significantly enhances performance in long-horizon goal-conditioned tasks.
Task category Hyperparameters Methods
Environment Type Size $n$ $k$ $\beta^{h}$ $\beta^{\ell}$ HIQL OTA
PointMaze navigate large 5 25 3.0 3.0 $58{\scriptstyle\pm 5}$ $\mathbf{85}{\scriptstyle\pm 5}$
giant 5 25 3.0 3.0 $46{\scriptstyle\pm 9}$ $\mathbf{72}{\scriptstyle\pm 6}$
stitch large 5 25 3.0 3.0 $13{\scriptstyle\pm 6}$ $\mathbf{46}{\scriptstyle\pm 7}$
giant 5 25 3.0 3.0 $0{\scriptstyle\pm 0}$ $\mathbf{44}{\scriptstyle\pm 8}$
AntMaze navigate large 5 25 3.0 3.0 $91{\scriptstyle\pm 2}$ $\mathbf{91}{\scriptstyle\pm 1}$
giant 5 25 3.0 3.0 $65{\scriptstyle\pm 5}$ $\mathbf{70}{\scriptstyle\pm 2}$
stitch large 5 25 3.0 3.0 $67{\scriptstyle\pm 5}$ $\mathbf{79}{\scriptstyle\pm 3}$
giant 5 25 3.0 3.0 $2{\scriptstyle\pm 2}$ $\mathbf{29}{\scriptstyle\pm 5}$
explore medium 5 25 3.0 3.0 $37{\scriptstyle\pm 10}$ $\mathbf{93}{\scriptstyle\pm 3}$
large 10 25 3.0 3.0 $4{\scriptstyle\pm 5}$ $\mathbf{62}{\scriptstyle\pm 12}$
HumanoidMaze navigate large 20 100 3.0 3.0 $49{\scriptstyle\pm 4}$ $\mathbf{82}{\scriptstyle\pm 2}$
giant 20 100 3.0 3.0 $12{\scriptstyle\pm 4}$ $\mathbf{91}{\scriptstyle\pm 1}$
stitch large 20 100 3.0 3.0 $28{\scriptstyle\pm 3}$ $\mathbf{43}{\scriptstyle\pm 3}$
giant 20 100 3.0 3.0 $3{\scriptstyle\pm 2}$ $\mathbf{61}{\scriptstyle\pm 3}$
Table 7: Performance under unified hyperparameters.
D.3 Performance on Visual Play Datasets
We evaluate the performance of OTA on visual play datasets, with results summarized in Table 8. For a fair comparison, we fix the hyperparameters $(k,\beta^{h},\beta^{l})=(10,3.0,3.0)$ for both HIQL and OTA, varying only $n=2$.
Task category Methods
Environment Data Env HIQL OTA
Visual-cube play double $48{\scriptstyle\pm 4}$ $\mathbf{51}{\scriptstyle\pm 3}$
play triple $21{\scriptstyle\pm 5}$ $\mathbf{28}{\scriptstyle\pm 1}$
Visual-scene play - $50{\scriptstyle\pm 5}$ $\mathbf{56}{\scriptstyle\pm 5}$
Table 8: Performance comparison on visual manipulation play dataset.